### v5

# Energy-and Spectral-Efficiency Trade-off in Distributed Massive-MIMO Networks

Abstract
This paper investigates a fundamental yet under-explored trade-off between energy efficiency (EE) and spectral efficiency (SE) in distributed massive MIMO (D-mMIMO) systems. Unlike conventional EE-SE trade-off studies that primarily focus on transmission power, D-mMIMO systems introduce new energy consumption factors—including fronthaul signaling and distributed signal processing—which are heavily influenced by AP-UE association. This work highlights the critical need for a system-level EE-SE trade-off framework that accounts for these unique aspects of D-mMIMO. We formulate a joint optimization problem that maximizes EE while satisfying uplink sum-SE constraints, through the coordinated design of power allocation and AP-UE association strategies. By explicitly considering both transmission and infrastructure-related energy costs, our approach enables energy-aware network design without compromising throughput. Numerical simulations demonstrate the substantial impact of dynamic AP-UE association and power control on the EE-SE trade-off, providing actionable insights for an efficient deployment of large-scale distributed MIMO networks in next-generation wireless systems.
Index Terms:
Distributed massive MIMO, energy efficiency, spectral efficiency, AP-UE association, power allocation, uplink communication, EE-SE trade-off.
I Introduction
The exponential growth in connected devices and the increasing demand for high data rates have placed unprecedented pressure on wireless networks, making both spectral efficiency (SE) and energy efficiency (EE) critical design considerations for 5G and beyond. While SE ensures higher throughput and better utilization of spectral resources, maximizing it often comes at the cost of increased power consumption. This leads to significant challenges in designing communication systems that are both high-performing and energy-efficient.
The distributed massive multiple-input multiple-output (D-mMIMO) systems have emerged as a transformative architecture for next-generation networks. By spatially distributing access points (APs) across a service area and jointly serving users without traditional cell boundaries, D-mMIMO systems improve macro-diversity, mitigate cell-edge issues, and enable uniform service quality [1, 2]. Within this architecture, two system-level factors—power allocation and access point-to-user equipment (AP-UE) association—play a central role in determining both EE and SE [1]. As the push for green communication and sustainable network design intensifies, it becomes imperative to explore and quantify the trade-off between EE and SE—a challenge that is particularly relevant in D-mMIMO systems.
Unlike centralized MIMO systems where transmit power dominates total energy consumption, D-mMIMO introduces significant additional energy costs due to distributed circuit operations, fronthaul communication, and cooperative signal processing [3]. In such systems, the total number of active APs and their associations with UEs have a more significant impact on system-wide energy consumption than the over-the-air transmit power. Therefore, evaluating the EE-SE trade-off in D-mMIMO systems from a system design perspective necessitates incorporating these architecture-specific energy components.
The EE-SE trade-off is explored in [4, 5, 6] for wireless MIMO systems, which predominantly focus on transmit power control. Although works related to distributed antennas systems, such as [7, 8, 9] consider additional power from backhaul links, they still restrict the optimization to transmit power, treating other power components as constants, thus failing to account for the dynamic nature of AP-UE associations and their implications on fronthaul and processing power.
In [10], the authors consider a duplex distributed MIMO system. For the uplink scenario, they assume that all APs serve all UEs, thereby limiting the optimization to transmit power alone. However, in D-mMIMO systems, the AP-UE association is equally critical. While increasing the number of APs serving a user typically improves SE, it can lead to diminishing EE returns due to heightened interference, coordination complexity, and increased system overhead. Hence, the number and spatial deployment of active APs, as well as their user associations, have a direct and significant impact on the network’s energy consumption. Therefore, the fixed-threshold-based or static association strategies are insufficient for balancing EE and SE in D-mMIMO deployments.
Some works have acknowledged AP-related power consumption, but those still fall short in addressing the EE-SE trade-off exhaustively. For example, the authors in [11] consider the AP-UE association and power allocation in downlink D-mMIMO for EE maximization, but without analyzing the EE-SE trade-off or implementing joint optimization. Similarly, [12] presents joint power control and active AP selection for downlink, yet assumes that all active APs serve all users, negating the benefit of dynamic AP-UE association.
Other studies have pursued SE maximization or power minimization but omit energy efficiency implications. In [13, 14, 15], uplink SE is optimized through separate AP selection and power allocation, while EE considerations are left out. Works like [16] address EE maximization but focus only on transmit power optimization. Recent studies such as [17, 18] consider joint AP-UE association and power allocation for SE maximization, but overlook energy consumption metrics. Even machine learning-based approach in [19], while proposing dynamic AP sleep modes, assumes a fixed AP-UE serving pattern and minimizes total energy consumption without explicitly optimizing EE, which may not align with energy-efficient operation.
A consistent limitation across much of the existing literature is the lack of dynamic, QoS-aware AP-UE association. Fixed or static serving strategies cannot adequately adapt to different user densities, SE requirements, or energy constraints. This gap is particularly impactful in large-scale deployments—such as smart cities, industrial IoT, or factory automation—where system-wide energy budgets and SE guarantees are simultaneously critical.
We address this critical gap by studying the joint impact of power allocation and the AP-UE association on the EE-SE trade-off in uplink D-mMIMO systems. We formulate an optimization problem that maximizes the overall energy efficiency while satisfying a minimum sum spectral efficiency requirement, thereby enforcing quality-of-service (QoS) constraints. Our approach enables the network to determine the optimal set of active APs and their corresponding UE assignments, adapting the serving configuration based on SE demands or EE priorities. This dynamic and system-aware resource allocation paradigm is essential for the sustainable design of future wireless networks.
The main contributions of this work are summarized as follows:
  • New Perspective on EE-SE Trade-off in D-mMIMO: We introduce a novel system-level framework for analyzing the trade-off between EE and SE in uplink D-mMIMO systems. Unlike conventional EE-SE analyses that focus solely on transmit power, our model incorporates realistic dynamic energy consumption from fronthaul signaling, and distributed signal processing, which are essential for understanding the energy dynamics of dense networks.
  • Joint Optimization of Power Allocation and AP-UE Association: We formulate and solve a joint optimization problem that simultaneously determines the transmit power levels and dynamic AP-UE association to maximize the overall EE of the system, subject to a sum-SE constraint. This joint design allows the system to dynamically adapt serving relationships based on SE requirements or EE priority, improving energy usage while ensuring QoS guarantees.
  • Deployment-Oriented Insights: This work offers practical guidance for D-mMIMO network deployment by identifying how the optimal number of APs depends on user density and SE requirements. It emphasizes the importance of dynamic AP-UE association and strategic scaling of AP–by turning APs ON or OFF based on EE-SE trade-off analysis–to achieve higher SE without incurring excessive energy costs from fronthaul and signal processing overhead.
  • Numerical Validation and Performance Evaluation: We validate the proposed framework through extensive numerical simulations, evaluating its performance under various system configurations. The results demonstrate the significant impact of AP density, user load, and SE requirements on the EE-SE trade-off, highlighting the importance of dynamic and joint resource allocation in energy-efficient D-mMIMO design.
Organization: The rest of the paper is structured as follows. Section II presents the system model. Section III describes the optimization problem. Section IV details the proposed solution methodology. Section V presents numerical results that analyze the EE-SE trade-off under various scenarios. Section VI concludes the paper and discusses future directions.
[…]
II System Model
We consider the uplink of a distributed massive MIMO system comprising $T$ single-antenna UEs and $M$ APs, each equipped with $A$ antennas, uniformly deployed over a defined coverage area. The total number of antennas in the system is thus $MA$, and we assume a large antenna-to-user ratio, i.e., $T\ll MA$, which enables significant spatial multiplexing gains and interference suppression capabilities.
Each AP is connected to a centralized processing unit (CPU) via reliable fronthaul link, enabling coordinated processing, data exchange, and user scheduling across the network. The system operates under a user-centric transmission paradigm, where each UE is served by a subset of geographically proximate APs selected based on large-scale fading metrics, as illustrated in Fig. 1. This approach enhances scalability and reduces fronthaul signaling load compared to fully connected cell-free architectures.
Figure 1: Illustration of a user-centric distributed massive MIMO system. Each oval indicates a UE and the APs serving it.
We adopt time-division duplexing (TDD) for channel reciprocity and efficient spectrum usage. The system employs a block fading model, where the channel remains constant over a coherence interval of $L_{c}$ symbols and changes independently across blocks. Each coherence block dedicates $L_{p}$ symbols for uplink pilot transmission, where all $T$ UEs transmit orthogonal pilot sequences for channel estimation. All APs operate over the same time-frequency resources and are capable of simultaneously serving multiple users.
Let $\mathbf{h}_{mt}\in\mathbb{C}^{A\times 1}$ denote the small-scale fading vector between the $m$-th AP and the $t$-th UE. We assume independent Rayleigh fading, such that $\mathbf{h}_{mt}\sim\mathcal{CN}(0,\mathbf{I}A)$, where all components are i.i.d. complex Gaussian with zero mean and unit variance. The large-scale fading coefficient (LSFC) $\beta{mt}$ accounts for path loss and shadowing. It is assumed to be constant over many coherence intervals and known at the network level. Thus, the overall channel vector from the $t$-th UE to the $m$-th AP is given by: $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$.
II-A Uplink Pilot Training
We consider uplink channel estimation under TDD operation, where each UE $t\in\{1,\dots,T\}$ transmits a pilot sequence $\sqrt{L_{p}}\bm{\psi}_{t}\in\mathbb{C}^{L_{p}\times 1}$, satisfying $\|\bm{\psi}_{t}\|^{2}=1$. The received pilot signal at AP $m\in\{1,\dots,M\}$, denoted by $\mathbf{Y}^{\mathrm{pilot}}_{m}\in\mathbb{C}^{A\times L_{p}}$, is expressed as:
$\mathbf{Y}^{\mathrm{pilot}}_{m}=\sum_{t=1}^{T}\sqrt{L_{p}p_{p}}\,\mathbf{g}_{mt}\bm{\psi}_{t}^{\mathsf{T}}+\mathbf{N}_{m},$
where $\mathbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ denotes the additive white Gaussian noise matrix with i.i.d. $\mathcal{CN}(0,\sigma^{2})$ entries, and $p_{p}$ is the uplink maximum pilot power.
To obtain the MMSE estimate $\hat{\mathbf{g}}_{mt}\in\mathbb{C}^{A\times 1}$ of the channel $\mathbf{g}_{mt}$, the AP $m$ correlates $\mathbf{Y}^{\mathrm{pilot}}_{m}$ with the conjugate of UE $t$’s pilot [1]:
$\hat{\mathbf{g}}_{mt}=\frac{\sqrt{L_{p}p_{p}}\beta_{mt}}{\sum_{t^{\prime}=1}^{T}L_{p}p_{p}\beta_{mt^{\prime}}|\bm{\psi}_{t}^{\mathsf{H}}\bm{\psi}_{t^{\prime}}|^{2}+\sigma^{2}}\,\mathbf{Y}^{\mathrm{pilot}}_{m}\bm{\psi}_{t}^{*}.$
The corresponding mean-squared value of the channel estimate is [1]:
$\gamma_{mt}=\mathbb{E}\{\|\hat{\mathbf{g}}_{mt}\|^{2}\}=\frac{L_{p}p_{p}\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}L_{p}p_{p}\beta_{mt^{\prime}}|\bm{\psi}_{t}^{\mathsf{H}}\bm{\psi}_{t^{\prime}}|^{2}+\sigma^{2}}.$
Note that $\gamma_{mt}$ is a measure of channel estimation quality and is significantly affected by pilot contamination (i.e., non-orthogonal pilot reuse).
II-B Uplink Data Transmission
During the data transmission phase, the received uplink signal at AP $m$ is:
$\mathbf{y}^{\mathrm{ul}}_{m}=\sqrt{p_{u}}\sum_{t=1}^{T}\mathbf{g}_{mt}\sqrt{\eta^{u}_{t}}x_{t}+\mathbf{n}_{m},$ (1)
where $x_{t}$ is the data symbol transmitted by the UE $t$, satisfying $\mathbb{E}\{|x_{t}|^{2}\}=1$; $\eta^{u}_{t}\in[0,1]$ is the uplink power control coefficient for the UE $t$; and $p_{u}$ denotes the uplink maximum power. The noise vector $\mathbf{n}_{m}\in\mathbb{C}^{A\times 1}$ contains i.i.d. $\mathcal{CN}(0,\sigma^{2})$ entries.
Each AP applies a local combining vector $\mathbf{v}_{mt}\in\mathbb{C}^{A\times 1}$ to detect the signal of the UE $t$. The partial detection $\hat{y}_{mt}=\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{y}^{\mathrm{ul}}_{m}$ is forwarded to the CPU. The CPU performs Large-Scale Fading Decoding (LSFD) across the APs [3]:
$\hat{y}_{t}=\sum_{m=1}^{M}d_{mt}a_{mt}\hat{y}_{mt}=\sum_{m=1}^{M}d_{mt}a_{mt}\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{y}^{\mathrm{ul}}_{m},$ (2)
where $d_{mt}\in\{0,1\}$ is the AP-UE association indicator: $d_{mt}=1$ if AP $m$ serves UE $t$; otherwise $d_{mt}=0$. Also, $a_{mt}$ is the LSFD coefficient for the UE $t$ with respect to the AP $m$.
Substituting $\mathbf{y}^{\mathrm{ul}}_{m}$ in (2), the overall estimated signal becomes [20]:
$\hat{y}_{t}=$ $\underbrace{\sum_{m=1}^{M}\sqrt{\eta^{u}_{t}p_{u}}\,d_{mt}a_{mt}\,\mathbb{E}\left\{\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{g}_{mt}\right\}x_{t}}_{\text{Desired signal}}+\underbrace{\sum_{m=1}^{M}\sqrt{\eta^{u}_{t}p_{u}}\,d_{mt}a_{mt}\left(\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{g}_{mt}-\mathbb{E}\left\{\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{g}_{mt}\right\}\right)x_{t}}_{\text{Beamforming uncertainty}}$
$+\underbrace{\sum_{k\in\mathcal{P}_{t}\setminus\{t\}}\sum_{m=1}^{M}\sqrt{\eta^{u}_{k}p_{u}}\,d_{mt}a_{mt}\,\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{g}_{mk}x_{k}}_{\text{Coherent interference (pilot sharing)}}+\underbrace{\sum_{k\notin\mathcal{P}_{t}}\sum_{m=1}^{M}\sqrt{\eta^{u}_{k}p_{u}}\,d_{mt}a_{mt}\,\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{g}_{mk}x_{k}}_{\text{Non-coherent interference}}$
$+\underbrace{\sum_{m=1}^{M}d_{mt}a_{mt}\mathbf{v}_{mt}^{\mathsf{H}}\mathbf{n}_{m}}_{\text{Noise}}.$ (3)
Here, $\mathcal{P}_{t}$ denotes the set of UEs sharing the same pilot sequence as the UE $t$. This decomposition separates the desired signal, various interference terms, and noise, providing a foundation for the achievable SE analysis in the next section.
II-C Spectral Efficiency and Energy Efficiency
We have considered Partial Full-Pilot Zero-Forcing (PFZF) combining [20], where each AP suppresses the interference caused by users with strong channels from other users who also have strong channels, using local zero-forcing constraints—provided the AP has sufficient spatial degrees of freedom. Specifically, each AP $m$ classifies the users it serves into two groups: the strong users, denoted by $\mathcal{S}_{m}$, whose interference is actively suppressed, and the weak users, denoted by $\mathcal{W}_{m}$, whose interference is only partially mitigated due to limited spatial resources. Correspondingly, for each user $t$, $\mathcal{M}_{t}$ represents the set of APs for which the user $t$ is considered strong (i.e., $t\in\mathcal{S}_{m}$), and $\mathcal{Q}_{t}$ denotes the set of APs for which the user $t$ is considered weak (i.e., $t\in\mathcal{W}_{m}$).
The uplink SE of user $t$ is lower bounded as [20]:
$\mathrm{SE}^{u}_{t}=w\log_{2}\left(1+\underbrace{\frac{\mathrm{DS}_{t}}{\mathrm{PC}_{t}+\mathrm{BU}_{t}+\mathrm{NI}_{t}+\mathrm{N}_{t}}}_{\Gamma_{t}}\right),$ (4)
where $w=\frac{\left(1-\frac{L_{p}}{L_{c}}\right)}{2}$ is the pre-log factor accounting for pilot overhead. This lower bound is obtained using the bounding technique in [21], which guarantees a rigorous and tractable expression for large-scale analysis.
The term $\mathrm{DS}_{t}$ in (4) represents the desired signal component, given by
$\mathrm{DS}_{t}={p_{u}\eta^{u}_{t}}\left|\sum_{m\in\mathcal{Z}_{t}}d_{mt}a^{*}_{mt}\gamma_{mt}+A\sum_{m\in\mathcal{Q}_{t}}d_{mt}a^{*}_{mt}\gamma_{mt}\right|^{2},$ (5)
The pilot contamination term $\mathrm{PC}_{t}$ accounts for coherent interference from users sharing the same pilot and is expressed as
$\mathrm{PC}_{t}=\sum_{\ \ t^{\prime}\in\mathcal{P}_{t}/\{t\}}\eta^{u}_{t^{\prime}}p_{u}\Bigg{|}\sum\limits_{\ \ m\in\mathcal{Z}_{t}}\frac{d_{mt}a^{*}_{mt}\gamma_{mt}\sqrt{\eta_{t^{\prime}}}\beta_{mt^{\prime}}}{\sqrt{\eta_{t}}\beta_{mt}}+\!A\sum\limits_{m\in\mathcal{Q}_{t}}\frac{d_{mt}a^{*}_{mt}\gamma_{mt}\sqrt{\eta_{t^{\prime}}}\beta_{mt^{\prime}}}{\sqrt{\eta_{t}}\beta_{mt}}\Bigg{|}^{2}.$ (6)
The term $\mathrm{BU}_{t}$ captures the beamforming uncertainty caused by the mismatch between the actual channel and its estimate:
$\mathrm{BU}_{t}=p_{u}\eta^{u}_{t}\left(\sum_{m\in\mathcal{Z}_{t}}\frac{d_{mt}^{2}|a^{*}_{mt}|^{2}\gamma_{mt}(\beta_{mt}-\gamma_{mt})}{A-L_{\mathcal{S}_{m}}}+A\sum_{m\in\mathcal{Q}_{t}}d_{mt}^{2}|a^{*}_{mt}|^{2}\gamma_{mt}\beta_{mt}\right).$ (7)
Non-coherent interference from other interfering users is represented by $\mathrm{NI}_{t}$, given by
$\mathrm{NI}_{t}=\sum_{t^{\prime}\neq t}p_{u}\eta^{u}_{t^{\prime}}\left(\sum_{m\in\mathcal{Z}_{t}}\frac{d_{mt}^{2}|a^{*}_{mt}|^{2}\gamma_{mt}(\beta_{mt^{\prime}}-\gamma_{mt^{\prime}})}{A-L_{\mathcal{S}_{m}}}+A\sum_{m\in\mathcal{Q}_{t}}d_{mt}^{2}|a^{*}_{mt}|^{2}\gamma_{mt}\beta_{mt^{\prime}}\right).$ (8)
Finally, $\mathrm{N}_{t}$ accounts for the thermal noise after combining:
$\mathrm{N}_{t}=\sum_{m\in\mathcal{Z}_{t}}\frac{d_{mt}^{2}|a^{*}_{mt}|^{2}\gamma_{mt}}{A-L_{\mathcal{S}_{m}}}+A\sum_{m\in\mathcal{Q}_{t}}d_{mt}^{2}|a^{*}_{mt}|^{2}\gamma_{mt}.$ (9)
The total uplink energy efficiency of the network (in bits per joule) is defined as:
$\text{EE}=\frac{wB\sum_{t=1}^{T}\log_{2}\left(1+\Gamma_{t}\right)}{P_{\text{T}}},$ (10)
where $B$ is the total bandwidth and, following [3],
$P_{\text{T}}$ $=P^{\text{fix}}_{\text{T}}+P^{c}_{\text{T}}+wP^{\text{deco}}_{cpu}B\sum_{t=1}^{T}\log_{2}\left(1+\Gamma_{t}\right),$
$P^{c}_{\text{T}}$ $=\sum_{t=1}^{T}\left(\frac{\eta^{u}_{t}p_{u}}{\zeta}+d_{mt}P^{\text{lsfd}}_{cpu}\right)+\sum_{m=1}^{M}\left(Ad_{mt}P^{\text{proc}}+d_{mt}P^{\text{sig}}\right),$
$P^{\text{fix}}_{\text{T}}$ $=TP^{\text{c}}_{ue}+MAP^{\text{c}}_{ap}+MP^{\text{fix}}_{fh}+P^{\text{fix}}_{cpu}.$
Here, $P^{\text{c}}_{ue}$, $P^{\text{c}}_{ap}$, $P^{\text{proc}}$, $P^{\text{fix}}_{fh}$, $P^{\text{sig}}$, $P^{\text{fix}}_{cpu}$, $P^{\text{lsfd}}_{cpu}$, and $P^{\text{deco}}_{cpu}$ denote the circuit power of UEs and APs, signal processing power, fronthaul power, and CPU processing powers for LSFD and decoding, respectively. The parameter $\zeta\in(0,1]$ denotes the power amplifier efficiency.
III Problem Formulation
In this section, we formulate an optimization problem aimed at maximizing the EE by jointly optimizing the power allocation for all UEs and the AP-UE associations. This is subject to meeting a minimum sum SE requirement. The optimization problem is expressed as:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+P^{\text{deco}}_{cpu}B\sum_{t=1}^{T}\text{SE}^{u}_{t}},$ (11a)
subject to: $d_{mt}\in\{0,1\},\quad\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (11b)
$0\leq\eta^{u}_{t}\leq 1,\quad\forall t\in\mathcal{T},$ (11c)
$w\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})\geq\text{SE}^{\text{QoS}},$ (11d)
$\sum\limits_{m=1}^{M}d_{mt}\geq 1,\quad\forall t\in\mathcal{T},$ (11e)
where $\bm{\eta^{u}}=\{\eta^{u}_{t}\}_{t\in\mathcal{T}}$ is the set of uplink power control coefficients, and D is the AP-UE association matrix whose element $d_{mt}$ is binary, indicating whether AP $m$ serves UE $t$. The term $\text{SE}^{\text{QoS}}$ denotes the minimum total SE required to satisfy QoS constraints.
Constraint (11b) enforces binary AP-UE association. Constraint (11c) ensures that the power control coefficients lie within a feasible range. Constraint (11d) ensures that the aggregated SE meets the QoS threshold, and (11e) guarantees that each UE is served by at least one AP.
The problem defined in (11) is a mixed-integer nonlinear programming (MINLP) problem. Such problems are generally NP-hard due to their combinatorial and non-convex nature. Therefore, to reduce the computational complexity, we adopt the fractional programming and quadratic transformation techniques as proposed in [22], as described in the next section.
IV Fractional Programming-Based Energy Efficiency Maximization
As discussed earlier, the optimization problem in (11) is computationally challenging due to its mixed-integer and non-convex nature. To obtain a feasible and tractable solution that satisfies the required system-wide sum SE, we propose a solution framework based on fractional programming and quadratic transformation techniques. The approach is composed of several key steps, as outlined below:
  1. 1.
    Reformulation: To reduce the computational cost while preserving optimality, we first reformulate the objective function in (11) to a more tractable form.
  2. 2.
    Binary Relaxation: The mixed-integer nature of the problem is handled by relaxing the binary AP-UE association variables $d_{mt}\in\{0,1\}$ to continuous values in the interval $[0,1]$, thereby converting the MINLP into a non-linear programming (NLP) problem.
  3. 3.
    Quadratic Transformation for SINR: The non-linear expression inside the logarithm, $\Gamma_{t}$, is made tractable by introducing auxiliary variables and applying the quadratic transformation method from [22]. This helps in approximating and linearizing the non-convex SINR term.
  4. 4.
    Handling Fractional Objective Function: The fractional structure of the EE objective is transformed into an equivalent, tractable form using Dinkelbach’s method or similar auxiliary variable-based techniques in conjunction with the quadratic transformation, as in [22].
  5. 5.
    Decoupling of Variables: The coupling between the power control vector $\bm{\eta^{u}}$ and the association matrix D introduces additional non-convexity. This is resolved using the decoupling strategy employed in [18], which transforms the problem into a convex one by alternating optimization or successive convex approximation.
  6. 6.
    Iterative Refinement: To provide accurate solutions that approximate the original problem, we employ an iterative algorithm that refines the auxiliary variables and updates the relaxed variables until convergence is achieved or a predefined tolerance is met.
This structured approach significantly reduces the complexity of solving the EE maximization problem (11) while preserving accuracy and feasibility with respect to the original constraints.
Step 1: Reformulating the Objective Function: In the original objective function (11a), both the numerator and the denominator contain the sum SE term. This increases the complexity of solving the problem. Therefore, without affecting the optimality of the solution, we reformulate the objective by removing the $P^{\text{deco}}_{cpu}wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})$ term, which is monotonically increasing with respect to the sum SE, from the denominator. Specifically, we consider a simplified energy efficiency expression where the denominator only includes the static and circuit power consumption terms. The modified optimization problem becomes:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (12a)
subject to: $d_{mt}\in\{0,1\},\forall m\in\mathcal{M},\ t\in\mathcal{T},$ (12b)
$0\leq\eta^{u}_{t}\leq 1,\forall t\in\mathcal{T},$ (12c)
$w\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})\geq\text{SE}^{\text{QoS}},$ (12d)
$\sum\limits_{m=1}^{M}d_{mt}\geq 1,\forall t\in\mathcal{T},$ (12e)
Since the removed term $P^{\text{deco}}_{cpu}wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})$ is monotonically increasing with respect to the sum SE, maximizing the modified objective function in (12a) still leads to the same optimal solution as the original formulation in (11a). Therefore, solving Problem (12) is equivalent to solving the original Problem (11), but with a lower computational cost.
Step 2: Relaxation of the Binary Association Variable: To mitigate the high computational complexity associated with solving mixed-integer non-linear programming (MINLP) problems, we relax the binary constraint on the AP-UE association variable $d_{mt}$. Specifically, the binary constraint in (12b) is relaxed to a continuous one as follows:
$0\leq d_{mt}\leq 1,\quad\forall m\in\mathcal{M},\ t\in\mathcal{T}.$ (13)
Remark 1: Although the variable $d_{mt}$ is now allowed to take continuous values, the structure of the objective function in (12a), particularly the power consumption term, implicitly penalizes fractional values. This is because the association cost increases with the number of active AP-UE links. As a result, the optimization naturally encourages $d_{mt}$ to converge towards binary values (either close to 0 or 1), thereby approximately satisfying the original binary constraint in (12b) while significantly reducing the problem’s complexity.
Step 3: Approximation of the SINR Expression via Auxiliary Variables and Quadratic Transformation: The SINR term $\Gamma_{t}$ in the objective function (12a) and constraint (12d) is inherently non-linear and non-convex due to its fractional structure. To facilitate tractable optimization, we introduce an auxiliary variable $\Gamma^{*}_{t}$, which serves as a concave lower bound to $\Gamma_{t}$. By replacing $\Gamma_{t}$ with $\Gamma^{*}_{t}$, we effectively remove the non-linearity within the logarithmic function. This substitution necessitates a new constraint to ensure the validity of the approximation:
$\Gamma^{*}_{t}\leq\Gamma_{t},\quad\forall t\in\mathcal{T}.$ (14)
However, constraint (14) remains non-convex due to the fractional structure of $\Gamma_{t}$. To address this issue, we apply the quadratic transformation technique proposed in [22], which introduces an iterative concave lower bound on $\Gamma_{t}$. Consequently, the non-convex constraint (14) is replaced by a tractable approximation:
$\Gamma^{*}_{t}\leq 2z_{t}\sqrt{\mathrm{DS}_{t}}-z_{t}^{2}I_{t}.$ (15)
where $I_{t}=\mathrm{PC}_{t}+\mathrm{BU}_{t}+\mathrm{NI}_{t}+\mathrm{N}_{t}$ represents the aggregate interference and noise affecting UE $t$. The feasibility of constraint (15) lies in the fact that:
$\Gamma_{t}=\frac{\mathrm{DS}_{t}}{I_{t}}\geq 2z_{t}\sqrt{\mathrm{DS}_{t}}-z_{t}^{2}I_{t},$ (16)
The auxiliary variable $z_{t}$ is updated in each iteration using the current estimate of $\Gamma_{t}$, ensuring the concave lower bound progressively tightens.
Lemma 1.
: The quadratic transformation in (16) provides a concave lower bound for $\Gamma_{t}$. Therefore, constraint (15) serves as an approximation of the original non-convex constraint (14), and its substitution preserves the feasibility of Problem (12). Furthermore, when $z_{t}=z^{*}_{t}$, the lower bound becomes tight and constraint (15) is equivalent to (14), where
$z^{*}_{t}=\frac{\sqrt{\mathrm{DS}_{t}}}{I_{t}}.$ (17)
Proof: The quadratic transformation $2z_{t}\sqrt{\mathrm{DS}_{t}}-z_{t}^{2}I_{t}$ is concave over the $z_{t}$. The maximum of this concave function occurs at $z_{t}=z^{*}_{t}$, as defined in (17). At this value, the quadratic transformation becomes equal to the original SINR expression $\Gamma_{t}$, thereby establishing it as a valid lower bound. Hence, using this transformation in place of the original constraint retains the feasibility of original problem.
Incorporating the auxiliary variable $\Gamma^{*}_{t}$ and its corresponding constraint (15), the optimization problem in (12) is reformulated as:
$\underset{\bm{\eta^{u}},\textbf{D},\bm{z},\bm{\Gamma^{*}}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (18a)
subject to: $w\sum_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})\geq\text{SE}^{\text{QoS}},\quad\forall t\in\mathcal{T},$ (18b)
[…]
where $\bm{z}=\{z_{t}\}_{t\in\mathcal{T}}$ and $\bm{\Gamma^{*}}=\{\Gamma^{*}_{t}\}_{t\in\mathcal{T}}$.
According to Lemma 1, as the iterative process refines $z_{t}$, the auxiliary variable $\Gamma^{*}_{t}$ asymptotically converges to the true SINR $\Gamma_{t}$, thus ensuring that Problem (18) yields a solution equivalent to that of the original problem in (12).
Step 4: Handling Fractional Objective Function: Since the objective function in (18a) has a fractional form, we introduce two auxiliary variables $u$ and $v$, representing the numerator and denominator, respectively, to facilitate a more tractable reformulation. The optimization problem is rewritten as:
$\underset{\bm{\eta^{u}},\textbf{D},u,v,\bm{z},\bm{\Gamma^{*}}}{\max~{}}$ $\ \ \frac{u}{v}$ (19a)
subject to: $w\sum_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})\geq\text{SE}^{\text{QoS}},\ \forall t\in\mathcal{T},$ (19b)
$wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})\geq u,$ (19c)
$P^{\text{fix}}_{\text{T}}+\bar{P}_{\text{T}}\leq v,$ (19d)
[…]
Solving Problem (19) is equivalent to solving Problem (18). Since the objective is to maximize the ratio $\frac{u}{v}$, the optimal solution occurs when the numerator and denominator constraints, in (19c) and (19d) respectively, are satisfied with equality.
However, the objective function (19a) still retains a fractional form. To handle this, we apply the quadratic transformation technique from [22] to derive a concave lower bound:
$\frac{u}{v}\geq 2b\sqrt{u}-b^{2}v,$ (20)
where $b$ is an auxiliary variable updated iteratively. This substitution converts the original non-convex objective into a concave function, allowing efficient optimization.
Lemma 2.
: The quadratic transformation $2b\sqrt{u}-b^{2}v$ provides a concave lower bound for the ratio $\frac{u}{v}$. Thus, the right-hand side of (20) serves as a tractable approximation of the original non-linear objective function in (18a), and replacing the objective with this expression preserves the feasibility of Problem (19). Furthermore, when $b=b^{*}$, this lower bound becomes tight and the inequality in (20) becomes an equality, where
$b^{*}=\frac{\sqrt{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})}}{P^{\text{fix}}_{\text{T}}+P^{c}_{\text{T}}}.$ (21)
Proof: The quadratic transformation $2b\sqrt{u}-b^{2}v$ is concave in the variable $b$. The maximum value of this transformation is attained at $b=b^{*}$, as defined in (21), which satisfies $2b^{*}\sqrt{u}-(b^{*})^{2}v=\frac{u}{v}$. Therefore, the expression $2b\sqrt{u}-b^{2}v$ is a valid concave lower bound for $\frac{u}{v}$, and this bound is tight at the optimal value of $b=b^{*}$. Hence, the transformed objective function provides an accurate and efficient surrogate for the original fractional objective while improving tractability for optimization.
By substituting the objective with the lower bound from (20), we arrive at the following reformulated problem:
$\underset{\bm{\eta^{u}},\textbf{D},u,v,b,\textbf{z},\bm{\Gamma^{*}}}{\max~{}}$ $\ \ 2b\sqrt{u}-b^{2}v,$ (22a)
[…]
As the iterations proceed, the value of the objective function in (22a) approaches that of the original fractional form in (19a), as established in Lemma 2. This implies that solving Problem (22) is equivalent to solving Problem (19).
Step 5: Decoupling of Variables: After fixing the parameters $b$ and z, the only remaining non-convex term in Problem (22) is constraint (15), due to the coupling between $\bm{\eta^{u}}$ and D. It is important to note that the right-hand side of constraint (15) is bi-concave in $\bm{\eta^{u}}$ and D. Therefore, Problem (22) can be efficiently solved using an alternating optimization approach.
We first fix D, $b$, and z, and optimize over $\bm{\eta^{u}}$, $u$, $v$, and $\bm{\Gamma^{*}}$ by solving the following problem:
$f_{1}=\underset{\bm{\eta^{u}},u,v,\bm{\Gamma^{*}}}{\max~{}}$ $\ \ 2b\sqrt{u}-b^{2}v,$ (23a)
[…]
Once Problem (23) is solved, we update the values of $b$ and z accordingly. Next, we fix $\bm{\eta^{u}}$, $b$, and z, and optimize over D, $u$, $v$, and $\bm{\Gamma^{*}}$ by solving the following problem:
$f_{2}=\underset{\textbf{D},u,v,\bm{\Gamma^{*}}}{\max~{}}$ $\ \ 2b\sqrt{u}-b^{2}v,$ (24a)
[…]
Step 6: Iterative Solution: After solving Problem (24), the auxiliary variables $b$ and z are updated, and Problem (23) is solved again. This alternating procedure between the two sub-problems continues until convergence. The complete iterative procedure is summarized in Algorithm 1, and their visual representation of steps involved for optimization is provided in Fig. 2.
1:  Initialization: Initialize feasible values ${\bm{\eta^{u}}}^{(0)}$, $\textbf{D}^{(0)}$, and set iteration index $i=0$, tolerance $\epsilon=5\times 10^{-3}$.
2:  repeat
3:     $i\leftarrow i+1$
4:     for all  $t\in\mathcal{T}$  do
5:        Update $z_{t}$ using (17).
6:     end for
7:     Update $b$ using (21).
8:     Solve Problem (23) to update ${\bm{\eta^{u}}}^{(i)}$.
9:     for all  $t\in\mathcal{T}$  do
10:        Recompute $z_{t}$ using (17).
11:     end for
12:     Recompute $b$ using (21).
13:     Solve Problem (24) to update $\textbf{D}^{(i)}$.
14:     Set ${\bm{\eta^{u}}}^{(i+1)}\leftarrow{\bm{\eta^{u}}}^{(i)}$, $\textbf{D}^{(i+1)}\leftarrow\textbf{D}^{(i)}$.
15:  until  $\left|\frac{f_{2}^{(i+1)}-f_{2}^{(i)}}{f_{2}^{(i)}}\right|\leq\epsilon$.
Algorithm 1 Proposed Alternating Optimization Algorithm
Start / System ParametersReformulate Objective Function :  $\frac{wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+P^{\text{deco}}_{cpu}wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}\xRightarrow{}\frac{wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}$elax Binary AP-UE Variables : $d_{mk}\in\{0,1\}\xRightarrow{}d_{mk}\in[0,1]$Apply Quadratic Transformation for SINR : $\frac{\mathrm{DS}_{t}}{I_{t}}\xRightarrow{}2z_{t}\sqrt{\mathrm{DS}_{t}}-z_{t}^{2}I_{t}$gain Apply Quadratic Transformation for EE Fraction : $\frac{u}{v}\xRightarrow{}2b\sqrt{u}-b^{2}v$Decouple Power Factor ($\bm{\eta^{u}}$) and Association Variables ($\mathbf{D}$) by Solving Optimization AlternativelyIterative Refinement and Convergence CheckConverged EE
Figure 2: Flowchart of the proposed optimization framework for EE-SE trade-off in D-mMIMO systems.
Convergence Analysis:
The convergence of the proposed alternating optimization algorithm (Algorithm 1) to a stationary point of the original problem (11) is established under the following assumptions:
  • The feasibility of the original problem (11) is preserved at each iteration of Algorithm 1.
  • The objective function exhibits monotonic improvement at every iteration.
  • The objective functions of sub-problems (23) and (24) are bounded from above.
Theorem 1.
: Let the sequence $\{{\bm{\eta^{u}}}^{(i)},\textbf{D}^{(i)}\}$ be generated by Algorithm 1. Then, under the above assumptions 1–3 the sequence $\{{\bm{\eta^{u}}}^{(i)},\bm{D}^{(i)}\}$ converges to a limit point $\{{\bm{\eta^{u}}}^{*},\textbf{D}^{*}\}$, which is a stationary point of the original problem (11).
Proof: See Appendix A.
Furthermore numerical evidence of the speed of convergence as well as robustness of algorithm are discussed in the next section.
V Numerical Simulations
In this section, we present numerical simulations to evaluate the performance of the proposed joint power allocation and AP-UE association algorithm to maximize the EE under the sum SE constraint. The results include an analysis of the achieved EE-SE trade-off under varying network conditions, as well as the speed of convergence and robustness of the proposed optimization framework.
We conduct numerical simulations over a $1\times 1$ km2 geographic area where APs and UEs are independently and identically distributed over a square area, centred at the origin. Specifically, the x and y coordinates of each UE and AP are independently drawn from a continuous uniform distribution ranging from -0.5 to 0.5. To emulate an infinitely large network and eliminate edge effects, we adopt a wrap-around topology as described in [15]. In our setup, the number of APs is denoted by $M$, and the number of UEs is $T$.
The large-scale fading coefficients are modeled using a three-slope path loss model, with shadow fading applied using a log-normal distribution with 8 dB standard deviation. Unless specified otherwise, we fix the system parameters to $B=20$MHz $A=8$, $L_{c}=200$, and $L_{p}=5$. During the channel estimation phase, all pilots are transmitted at the maximum power of 100 mW. The initial UE transmit power for data transmission is also set to 100 mW. For the initial AP-UE association, each AP serves only those UEs for which it contributes at least 95% of the maximum large-scale fading coefficient (LSFC) across all APs. Also, $P^{\text{c}}_{ue}=100$ mW , $P^{\text{c}}_{ap}=100$ mW, $P^{\text{proc}}=800$, mW $P^{\text{fix}}_{fh}=825$ mW, $P^{\text{sig}}=10$ mW, $P^{\text{fix}}_{cpu}=5000$ mW, $P^{\text{lsfd}}_{cpu}=1000$, and $P^{\text{deco}}_{cpu}=1000$ mW/Gb/s [3]. Unless stated otherwise, all other parameters follow the configuration described in [1]. All numerical results are averaged over 50 independent simulation realizations.
V-A Performance Evaluation
Figure 3: The EE versus the sum SE threshold versus number of APs for $K=30$.
Fig. 3 presents a 3D plot showing the variation of energy efficiency (EE) with respect to the sum spectral efficiency (SE) threshold $\text{SE}^{\text{QoS}}$ and the number of APs $M$. For a fixed number of APs, increasing the SE requirement results in a consistent decline in EE. Specifically, for $M=30$, EE drops sharply from 3.6 to 0.2 Mbit/Joule as $\text{SE}^{\text{QoS}}$ increases from 55 to 75 bit/s/Hz, and becomes zero beyond 80 bit/s/Hz, indicating infeasibility. Similarly, for $M=60$, EE decreases from 4.22 to 0.3 Mbit/Joule as the SE threshold increases from 55 to 95 bit/s/Hz, and drops to zero for higher thresholds. For $M=110$, the EE decreases gradually from 3.8 to 0.5 Mbit/Joule as $\text{SE}^{QoS}$ rises from 55 to 110 bit/s/Hz. This trend is consistent across all $M$, demonstrating that higher SE targets lead to increased power consumption and hence reduced EE.
The degradation in EE with increasing $\text{SE}^{\text{QoS}}$ is attributed to the need for more aggressive AP-UE associations and increased resource allocation to meet the higher throughput demands. This leads to a rise in total power consumption, including AP circuit power, fronthaul transmission power, signal processing power, and CPU computation power. Thus, indicating the trade-off between the EE and SE of the deployed network. Importantly, the systems with a larger number of APs are able to support higher $\text{SE}^{\text{QoS}}$ demands, due to improved spatial diversity and better aggregate channel conditions. In contrast, network with fewer APs become infeasible at high SE thresholds due to insufficient coverage or capacity.
Furthermore, for a fixed $\text{SE}^{\text{QoS}}$, we observe a non-monotonic behavior of EE with respect to $M$. Initially, EE improves with increasing $M$ due to enhanced spatial diversity and improved channel conditions, which increase SE faster than the corresponding increase in power consumption. However, beyond a certain point, the power overhead associated with deploying more APs—especially circuit and processing power—starts to dominate, leading to a decline in EE. This trade-off highlights the importance of optimal AP density for ensuring energy-efficient operation while meeting QoS requirements in practical system deployments.
Figure 4: The EE versus the sum SE threshold versus number of UEs for $p_{n}p_{u}=0.1$.
Fig. 4 presents a 3D plot illustrating the EE as a function of the sum SE threshold $\text{SE}^{\text{QoS}}$ and the number of UEs $T$. For a fixed number of UEs, increasing the SE requirement consistently leads to a decline in EE. Specifically, for $T=20$, EE drops sharply from 2.9 to 0.3 Mbit/Joule as $\text{SE}^{\text{QoS}}$ increases from 55 to 65 bit/s/Hz and becomes zero beyond 70 bit/s/Hz, indicating that this setup cannot support higher SE demands. Similarly, for $T=40$, EE decreases from 3.9 to 0.16 Mbit/Joule as the SE threshold increases from 55 to 100 bit/s/Hz and becomes zero thereafter. For $T=80$, the EE gradually declines from 2.6 to 0.2 Mbit/Joule as $\text{SE}^{\text{QoS}}$ increases from 55 to 110 bit/s/Hz. This pattern is consistent across all UE counts, demonstrating that higher SE requirements lead to greater power consumption and, consequently, lower EE.
This EE–SE trade-off emerges due to the system’s need to increase AP-UE associations and allocate more transmission and processing resources to meet the SE demands. Notably, the systems with a larger number of UEs can support higher $\text{SE}^{\text{QoS}}$ thresholds due to increased spectrum utilization.
Furthermore, for a fixed SE requirement, EE exhibits a non-monotonic relationship with the number of UEs. Initially, EE improves with increasing $T$ because the system benefits from multi-user diversity and improved spectrum utilization. However, beyond a certain point, inter-user interference and the overhead of supporting more users—including circuit power and processing power—begin to outweigh the benefits, causing EE to decline. Compared to the case of increasing AP density, increasing UE density leads to a more pronounced drop in EE due to stronger inter-user interference. This trade-off highlights the importance of carefully selecting the UE density to balance EE and QoS satisfaction in real-world deployments.
Figure 5: The EE versus the sum SE threshold versus the maximum uplink power.
Fig. 5 represents the 3-D plot between the EE, the sum SE threshold and the maximum uplink power ($p_{u}$) that UEs can attain. In this plot, the sum SE threshold varies from $40$ to $110$ bits/s/Hz, and the maximum uplink UE power is varied from $10$ to $500$ mW. For the maximum power level of 10 mW, the EE decreases from $3.3$ to $0.04$ Mbits/Joule as the $\text{SE}^{\text{QoS}}$ increases from $40$ to $70$ bits/s/Hz and beyound that the EE remains zero. For 50 mW, the EE reduces from $3.9$ to $0.08$ Mbits/Joule as the $\text{SE}^{\text{QoS}}$ increases from $40$ to $80$ bits/s/Hz and beyound that EE remains zero. A similar trend is observed at a power level of 200 mW, the EE reduces from $4.1$ to $0.09$ Mbits/Joule as the $\text{SE}^{\text{QoS}}$ increases from $40$ to $90$ bits/s/Hz and beyound that the EE remains zero. This decreasing pattern persists across all higher power levels, consistently showing a reduction in EE as $\text{SE}^{\text{QoS}}$ increases. This decline is attributed to the necessity of either allocating more APs to UEs, increasing the power allocated to UEs, or both, to enhance the SE. Consequently, this increases the overall power consumption and diminishes the network’s EE. Also, the higher power level can support higher SE thresholds. We note an increase in the EE with rising maximum uplink power up to a certain point for the fixed sum SE threshold. Notably, the EE begins to saturate beyond a power level of 200 W. The optimization algorithm, thereby, increases power consumption only when it contributes to the enhanced EE.
Figure 6: The EE versus the number of APs versus the number of UEs, when $\text{SE}^{\text{QoS}}=70$.
Fig. 6 presents the EE performance as a function of the number of APs and UEs under a moderate uplink sum SE threshold of $\text{SE}^{\text{QoS}}=70$ bit/s/Hz. The results highlight that, for a given user density, EE initially increases with the number of APs due to improved spatial diversity and reduced per-link transmit power. However, beyond an optimal point, EE begins to decline as the static power consumption associated with additional APs due to fronthaul signaling and circuit power outweighs the SE gains. For instance, when the network serves 30 UEs, the EE peaks at approximately 4.216 Mbits/Joule with 60 APs, after which it declines. Similarly, for 50 and 70 UEs, maximum EE values of 3.949 and 3.679 Mbits/Joule are attained at 90 and 100 APs, respectively. These trends underscore the importance of dynamically adapting AP activity to match user density. Specifically, in scenarios with fewer UEs, activating too many APs results in excessive overhead, while in denser user scenarios, a larger number of APs is necessary to maintain efficiency. Consequently, implementing AP sleep control mechanisms becomes essential for maximizing EE in practical deployments of D-mMIMO networks
Figure 7: The EE versus the number of APs versus the number of UEs, when $\text{SE}^{\text{QoS}}=100$.
Fig. 7 extends the EE analysis under a more stringent spectral efficiency constraint of $\text{SE}^{\text{QoS}}=100$ bit/s/Hz, offering comparative insights relative to Fig. 6. As expected, the overall EE values decrease to meet the higher QoS target. However, the optimal AP count—where EE is maximized—remains largely consistent for moderate to high user densities. For instance, with 50 and 70 UEs, the EE peaks remain at approximately 3.949 and 3.679 Mbits/Joule, attained with 90 and 100 APs respectively—matching the optimal AP counts observed in Fig. 6. In contrast, for a lower user load of 30 UEs, the EE maximum shifts to lower value of 3.633 Mbits/Joule at 120 APs, indicating that satisfying higher SE requirements in sparse user scenarios demands significantly more AP resources. These findings highlight the necessity of context-aware infrastructure adaptation, wherein AP activation should be dynamically tailored not only to user density but also to QoS targets, enabling scalable and energy-efficient D-mMIMO operation.
V-B Convergence Behaviour
Figure 8: Convergence behavior of the proposed algorithm in terms of EE versus number of iterations.
Fig. 8 shows the energy efficiency (EE) versus the number of iterations. Starting from an initial EE of approximately 0.6 Mbit/Joule, the proposed algorithm converges to around 2.7 Mbit/Joule within 3–5 iterations, with negligible improvement afterward. This indicates rapid convergence, making it suitable for dynamic network environments.
V-C Impact of Initialization
Figure 9: Final EE under different AP-UE association initializations for $M=40$ and $M=50$.
Fig. 9 illustrates the final EE after convergence under five different AP-UE association initialization schemes. For $M=40$, the EE varies within a narrow range of 4.02–4.06 Mbit/Joule, while for $M=50$, it lies between 4.11–4.17 Mbit/Joule.
The five initialization schemes are described as follows:
  • Scheme 1: Each UE is served by the top 10 APs based on the LSFC.
  • Scheme 2: Each AP serves its top 10 UEs based on LSFC.
  • Scheme 3: All APs serve all UEs (fully connected network).
  • Scheme 4: Each AP serves only those UEs for which it contributes at least 95% of the maximum LSFC across APs.
  • Scheme 5: Random AP-UE association.
Despite the diversity in initial association strategies, the resulting EE values are very close across all cases. This observation demonstrates the robustness of the proposed algorithm with respect to different AP-UE initializations.
Figure 10: Final EE under different UE transmit power initializations.
Fig. 10 presents the final energy efficiency (EE) achieved under five different UE transmit power initialization schemes. The final EE values converge within a narrow range of 4.00–4.03 Mbit/Joule. Thus, highlighting that the initialization of UE transmit power has a minimal effect on the converged solution. This consistency again validates the robustness of the proposed algorithm against different power initialization strategies.
VI Conclusion and Future Work
This work has systematically investigated the intricate trade-off between EE and SE in uplink distributed massive MIMO systems, accounting for both transmission and infrastructure-related power consumption. By jointly optimizing power allocation and AP-UE association, we have demonstrated how network configurations, particularly the number of active APs, can be dynamically adjusted to strike a balance between maximizing EE and meeting SE requirements. Notably, our results show that just increasing the number of APs to improve SE may lead to suboptimal EE performance due to elevated static and coordination power overheads.
From a system design perspective, the proposed framework provides essential guidance for deploying scalable and sustainable D-mMIMO networks. It enables network operators to adaptively activate or deactivate APs based on user density and SE demands, thereby avoiding energy over-provisioning and enhancing operational efficiency. Moreover, the AP-UE association plays a critical role in balancing fronthaul signaling power, further influencing overall energy consumption and coordination overhead. As future wireless networks such as 5G-Advanced and 6G increasingly prioritize high throughput and low latency, the insights from this work may be useful in highlighting the critical importance of integrating EE considerations into the core of radio resource management and infrastructure planning. Ultimately, this work supports the design of green and economically viable wireless communication systems without compromising quality of service.
A compelling future direction would involve drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO. Additionally, the influence of modulation and coding schemes, which is not considered in this study, presents an interesting direction for further exploration. Furthermore, incorporating dynamic traffic patterns could provide deeper insights into real-time energy-efficient network control.
Appendix A Proof of Theorem 1
  • Feasibility Preservation: By Lemma 2, the quadratic transformation $2b^{*}\sqrt{u}-(b^{*})^{2}v$ yields a tight concave lower bound for the fractional objective $\frac{u}{v}$, which becomes exact at $b=b^{*}$. This ensures equivalence between the transformed problem (22) and problem (18). Moreover, the quadratic transformation of the SINR term $\Gamma_{t}$ (as established in Lemma 1) ensures that the surrogate constraint (15) lower bounds $\Gamma_{t}$, thereby satisfying the original SINR constraint (12d). Thus, under the setting $z_{t}=z_{t}^{*}$, and following Remark 1, solving the approximated problem (18) guarantees equivalence with the original problem (11).
  • Monotonic Improvement: To demonstrate the non-decreasing nature of the optimization function $f(\bm{\eta^{u}},\textbf{D})=2b\sqrt{u}-b^{2}v$, assume that ${\bm{\eta^{u}}}^{*}$ represents the optimal value of $f$, when D is fixed. Given this, the inequality $f({\bm{\eta^{u}}}^{*},{\textbf{D}}^{(i)})\geq f({\bm{\eta^{u}}}^{(i)},{\textbf{D}}^{(i)})$ always holds due to the concavity of the function $f$ with respect to $\bm{\eta^{u}}$. When optimizing D to $\textbf{D}^{*}$, with $\bm{\eta^{u}}$ fixed at ${\bm{\eta^{u}}}^{*}$, the inequality $f({\bm{\eta^{u}}}^{*},{\textbf{D}}^{*})\geq f({\bm{\eta^{u}}}^{*},{\textbf{D}}^{(i)})$ always holds as $f$ is concave with respect to D. Therefore, combining these observations, we see that $f({\bm{\eta^{u}}}^{(i+1)},{\textbf{D}}^{(i+1)})\geq f({\bm{\eta^{u}}}^{(i)},{\textbf{D}}^{(i)})$, indicating $f$ is non-decreasing at each iteration. This non-decreasing trend makes the optimization function monotonically increasing in each iteration and also the optimization function is bounded from above, ensuring the convergence of the optimization algorithm (Algorithm 1), as the function does not increase indefinitely but plateaus at the maximum value.
  • Stationary Point: From the monotonic improvement and feasibility preservation, solving the transformed problem (22) via alternating optimization of subproblems (23) and (24) leads to a stationary point $\{{\bm{\eta^{u}}}^{*},\mathbf{D}^{*}\}$. Given the equivalence of the transformed problem to the original problem (discussed earlier), this stationary point also satisfies the original problem (11).